OpenStudy (anonymous):

The ratio of men to women working for a company 3 to 2. If there are 58 women working for the company, what is the total number of employees?

OpenStudy (stamp):

Consider your units involved. Remember that a ratio is a fraction, meaning that\[\frac{men}{women}*women=men\]

OpenStudy (kropot72):

To find the fraction of women in the company consider 5 employees in the ratio 3 men to 2 women. The fraction of women in the group is 2/5 and this fraction applies to the whole company. Therefore 58 women are 2/5 of the employees and 58/2 is 1/5 of the employees. To find the total number of employees multiply 58/2 by 5. \[total\ number\ of\ employees=\frac{58}{2}\times 5=you\ can\ calculate\]

OpenStudy (kropot72):

@stamp Are you there?

OpenStudy (stamp):

@kropot72 Yes

OpenStudy (kropot72):

Have you calculated the answer yet? Let me know if you need further explanation.

OpenStudy (stamp):

I do not need explanation. I can double check your answer if you would like confirmation

OpenStudy (kropot72):

@stamp Sorry. I meant to address my comments to the questioner (@rynewolfe).

OpenStudy (stamp):

\[women=58\]\[\frac{3\ men}{2\ women}58\ women=\frac{3(58)}{2}=3(29)=87\ men\]\[men+women=87+58=145\ employees\]